How do you estimate fractions? I have tried and tried and keep getting the estimate wrong. I get the actual answer right, but the estimate wrong. Here is an example
3 2/5*2 1/7 estimate is 35 and the actual answer is 32 31/40.
Thank you

Answer by Alan3354(69443)

How do you estimate fractions? I have tried and tried and keep getting the estimate wrong. I get the actual answer right, but the estimate wrong. Here is an example
3 2/5*2 1/7 estimate is 35 and the actual answer is 32 31/40.
--------------
3 2/5*2 1/7 = (17/5)*(15/7) = 51/7 = 7 2/7.
32 31/40 ???
------------------
To estimate, ir's ~ 3.5*2 = 7
----------------
What's the purpose of estimating?
A waste of time, IMO.